If two plane mirror are inclined at an angle of 450' calculate the number of images formed.
1
Expert's answer
2020-06-24T11:26:19-0400

As per the question,

The angle between two plane mirror is "\theta=45^\circ"

So, total number of image formed on the mirror, if object is not in the exactly middle of the two mirror, "n=\frac{360}{\theta}"

"n=\frac{360}{45}=8"

But if object is placed at the middle of the two mirror,

"n=\frac{360}{\theta}-1=8-1=7"